OpenStudy (anonymous):

Find the derivative of ln of absolute value of cos of pi/x. The answer is (pi/x^2)*tan(pi/x).

zepdrix (zepdrix):

\[\Large\bf\sf \ln\left|\cos \frac{\pi}{x}\right|\] Mmm ok absolute can be a bit of a doozy. Do you remember this derivative? \[\Large\bf\sf \frac{d}{dx}|x|\]It'll help when we apply one of our chains.

OpenStudy (anonymous):

No. Can you show the work?

zepdrix (zepdrix):

Sure, if we need to show the work for that.. we would do it like this:\[\Large\bf\sf |x|\quad:=\quad \sqrt{x^2}\]We'll take the derivative from this form.\[\Large\bf\sf \frac{d}{dx}\sqrt{x^2}\quad=\quad \frac{2x}{2\sqrt{x^2}}\quad=\quad \frac{x}{\sqrt{x^2}}\quad=\quad \frac{x}{|x|}\] \[\Large\bf\sf \implies \frac{d}{dx}|x|\quad=\quad \frac{x}{|x|}\]

zepdrix (zepdrix):

So there is a certain part in our problem when we chain rule, we'll have to take care of:\[\Large\bf\sf \frac{d}{dx}|\cos \frac{\pi}{x}|\]And we'll use this derivative rule.

OpenStudy (anonymous):

So it's cos pi/x divided by absolute value of cos pi/x?

zepdrix (zepdrix):

For that particular part, yes good. We would have to chain again though right? Because of the pi/x inside.

zepdrix (zepdrix):

\[\Large\bf\sf \color{royalblue}{\left(\ln\left|\cos \frac{\pi}{x}\right|\right)'}\quad=\quad \frac{1}{\left|\cos \dfrac{\pi}{x}\right|}\cdot \color{royalblue}{\left(\left|\cos\dfrac{\pi}{x}\right|\right)'}\]

zepdrix (zepdrix):

First step make sense? Log gives us one over the contents of the log right? Then we chain.

OpenStudy (anonymous):

Okay, how do we find the derivative of the inside?

zepdrix (zepdrix):

The inside is where we apply our fancy rule from before:\[\Large\bf\sf \frac{d}{dx}\left|\cos \frac{\pi}{x}\right|\quad=\quad \left(\left|\cos \frac{\pi}{x}\right|\right)'\]

zepdrix (zepdrix):

\[\Large\bf\sf =\frac{1}{\left|\cos \dfrac{\pi}{x}\right|}\cdot \frac{\cos\dfrac{\pi}{x}}{\left|\cos\dfrac{\pi}{x}\right|}\color{royalblue}{\left(\frac{\pi}{x}\right)'}\]

zepdrix (zepdrix):

Understand the second step? :o Took the derivative of absolute cosine, then chained again.

OpenStudy (anonymous):

Yes. Let me work it out. Please don't leave.

zepdrix (zepdrix):

k

zepdrix (zepdrix):

Oh crap I skipped a step, sorry about that. Our next inner chain is the stuff INSIDE of the absolute bars, not the stuff inside of cosine. I jumped ahead too far.

zepdrix (zepdrix):

\[\Large\bf\sf =\frac{1}{\left|\cos \dfrac{\pi}{x}\right|}\cdot \frac{\cos\dfrac{\pi}{x}}{\left|\cos\dfrac{\pi}{x}\right|}\color{royalblue}{\left(\cos\frac{\pi}{x}\right)'}\]

OpenStudy (anonymous):

I got cos pi/x divided by absolute value of cos pi/x squared times -sin pi/x and times -pi/x^2, how do I simplify this?

zepdrix (zepdrix):

\[\Large\bf\sf =\quad \frac{1}{\cos^2\dfrac{\pi}{x}}\cdot \left(\sin\frac{\pi}{x}\cos\frac{\pi}{x}\right)\cdot\frac{\pi}{x^2}\]Ok good. Cancel the negatives. Then you can DROP the absolute bars sinse the inside is a square. (Squares are always positive, so the absolute bars aren't doing anything for us anymore.

zepdrix (zepdrix):

See any nice cancellations or simplifications from there? :o

OpenStudy (anonymous):

Got it. Thanks. But is there another easier way to do this?

zepdrix (zepdrix):

If the absolute thing is confusing, you can just remember the definition of absolute and rewrite your problem as:\[\Large\bf\sf \frac{d}{dx}\ln \sqrt{\cos^2\frac{\pi}{x}}\]Because:\[\Large\bf\sf |x|\quad:=\quad \sqrt{x^2}\]And use normal rules of differentiation to do it from there. It's still a ton of chains, but it doesn't deal with the absolute anymore. This one is actually worth memorizing I think, it shows up later in Calc 2 and stuff.\[\Large\bf\sf \int\limits \tan x\;dx \quad=\quad -\ln|\cos x|\]\[\Large\bf\sf \implies \frac{d}{dx}\ln|\cos x|\quad=\quad -\tan x\]

OpenStudy (anonymous):

Thanks for the extra info!